### v5

@@ 6 Related Work
Requirements traceability (RT) has been extensively studied in RE mucha2024systematic ; li2023applications ; tufail2017systematic ; wang2018requirements ; ramesh1998factors . Existing work applies different technologies, ranging from traditional methods such as Information Retrieval (IR) and statistical models to more advanced approaches like Machine Learning (ML) and Deep Learning (DL). Early works borrowed IR techniques such as Vector Space Models (VSM), Latent Dirichlet Allocation (LDA), to find trace links between software artifacts via text relevancy gao2022using ; guo2017tackling ; kuang2017analyzing ; nishikawa2015recovering ; panichella2013and ; Mahmoud2013 ; Wang20193 ; sun2017frlink ; capobianco2013improving ; chhabra2017filtering ; bavota2014enhancing ; shao2013improved ; Wang20191 ; Li2020 ; moran2020improving . More advanced techniques have been introduced using ML Rasiman2022 ; Bella2019 ; mills2019tracing ; Bella2018 ; Mills2018 ; Rath2018 ; Zhao2018 ; Mills2017 ; Mills20172 ; Falessi2017 ; Hayes2015 ; le2015rclinker ; li2015recovering and DL kenton2019bert ; Zhang2021 ; Wang20192 ; Guo:17 ; Chen2019 ; Alhoshan2019 ; Alhoshan2018 ; Li2018 ; zhao2020extended ; Sultanov2013 , employing various algorithms — from classifiers like SVM, random forest, and decision trees to more sophisticated language models like BERT devlin-etal-2019-bert to find trace links. In recent years, with the emergence of LLMs, researchers have leveraged pre-trained knowledge through prompt engineering techniques to identify trace links between software artifacts hassine2024llm ; rodriguez2023prompts ; vogelsang2025impact . Hassine hassine2024llm proposed an LLM-based technique that uses zero-shot learning on GPT3.5 to find trace links between requirements and goals in Goal-oriented Language (GRL) models. Moreover, Rodriguez et al. rodriguez2023prompts proposed an approach that integrates zero-shot prompting with reasoning to enhance results in the Traceability Link Recovery (TLR) problem on diverse software artifacts. They have shown that a prompt that performs well with one model or dataset may not yield optimal results with another, underscoring the need to tailor prompts to the specific context. Recent studies have explored the use of prompting techniques for traceability within requirements fuchss2025lissa ; ronanki2024requirements ; fuchss2025beyond ; hey2025requirements . Hey et al. hey2025requirements used Retrieval-Augmented Generation (RAG) with LLMs for inter-requirements traceability, where the model first retrieves relevant candidate requirements. Then, it identifies trace links among the retrieved candidates using the KISS prompt, which consists of a yes/no question prompting the model to determine whether a trace link exists between a given pair of artifacts. Fuchß fuchss2025beyond et al. proposed an ensemble of LLM-based prompts for candidate filtering, where multiple LLM models are iteratively used to re-rank or filter the initial set of retrieved candidates. The prompt template used in this work follows the same design as that proposed by Hey et al. hey2025requirements . Additionally, Ronanki et al. ronanki2024requirements developed and experimented with five distinct interactive prompting patterns, each tuned for a specific dataset.
@@
Legal requirements traceability has only been investigated to a limited extent in the literature. Cleland-Huang et al. cleland:2010 propose a probabilistic approach that identifies trace links between requirements and the HIPAA regulation by computing probability values based on detecting indicator terms for regulations that correspond to requirements. The authors further propose extending the indicator terms with more domain-specific terms retrieved from the web. In a follow-up work, Gibiec et al. Gibiec:2010 further investigate mining the web. Guo et al. Guo:17 extend the previous two papers to improve the terminology gap problem, i.e., the mismatch between terms in requirements and regulations. The authors investigate different methods based on classification, ontologies, and web-mining and evaluate their approaches on HIPAA.
While previous research has made significant strides in requirements traceability using traditional IR methods and ML/DL techniques, these approaches exhibit notable limitations in addressing the complexities of the LRT task. Most notably, existing methods struggle with the terminology gap between regulations and technical requirements, do not generalize well across regulations, and lack adaptability to multi-domain applications. Additionally, studies that have proposed prompt-based approaches have not fully leveraged the capabilities of prompt engineering, providing only detailed instructions and limited context. Their approaches primarily rely on pre- and post-processing strategies hey2025requirements ; fuchss2025beyond , which in some cases query several large language models. In comparison to the above works, we empirically evaluate two automated LRT approaches: (1) a classifier-based solution leveraging sentence transformers and (2) a generative LLM-based solution guided by structured prompt engineering. By exploring these methods across two distinct regulations, HIPAA and GDPR, we advance the understanding of how modern NLP techniques can be adapted to meet the challenges of LRT. We also shed light on the possibilities, or lack thereof, of transfer learning across regulations. To the best of our knowledge, we are also among the first to identify the strengths and limitations of LLMs in this context. Further and larger studies with human experts are required to establish the benefits of LLMs for LRT.
@@ 7 Conclusion
This study presents a comparative evaluation of two approaches for Legal Requirements Traceability (LRT): a classifier-based method, $\mathcal{K}$ashif, leveraging sentence transformers, and a generative LLM-based method, Rice_LRT, designed using a structured prompt engineering framework. Moreover, we compare $\mathcal{K}$ashif against seven distinct baselines—ranging from traditional IR techniques to recent deep learning and transformer-based models (LSI, LDA, GloVe, TraceBERT, RoBERTa, and LLaMa2)—to comprehensively demonstrate its capabilities. Our results demonstrate that $\mathcal{K}$ashif provides significant improvements over the baselines in terms of F2 score, achieving a F2 score of $\approx$63% on HIPAA data (23 pp more than the best baseline). On the other hand, $\mathcal{K}$ashif’s performance deteriorates on more complex datasets such as GDPR, yielding only 15% recall. However, this also shows that $\mathcal{K}$ashif, despite being tuned on a similar but less complex dataset, achieves performance comparable to a simple prompt (answering yes/no) on GPT-4o, which does not fully exploit its capabilities.
Conversely, the Rice_LRT approach, built on generative LLMs, outperformed $\mathcal{K}$ashif on GDPR data, achieving an F2 score of 61% and reducing manual effort for traceability by enabling analysts to vet only a fraction of trace links. Moreover, Rice_LRT performed better than a simpler prompt, increasing TPs while reducing FPs. These findings suggest that generative LLMs and carefully designed prompts offer a promising approach to automating LRT tasks in complex legal domains. However, the approach has its challenges, including the occurrence of false positives that require further investigation. In addition to evaluating current state-of-the-art methods, this work highlights critical challenges, including terminology gaps between requirements and provisions By addressing these challenges, our study underscores the importance of tailoring solutions to the nuances of legal and regulatory contexts.
@@
Acknowledgements
The experiments conducted in this work were enabled by support provided by the Digital Research Alliance of Canada.
Declarations
Authors Contributions Romina Etezadi: Conceptualization, Investigation, Formal analysis, Software, Writing - original draft. Sallam Abualhaija: Conceptualization, Investigation, Data Curation, Formal analysis, Writing - original draft & review & editing. Chetan Arora: Conceptualization, Investigation, Data Curation, Formal analysis, Writing - review & editing. Lionel Briand: Conceptualization, Investigation, Formal analysis, Supervision, Writing - review & editing, Funding acquisition.
Funding This work is supported by the H2020 COSMOS European project, grant agreement No. 957254, NSERC of Canada under the Discovery and CRC programs, the Research Ireland grant 13/RC/2094-2, and the Luxembourg National Research Fund under grant number C23/IS/17958091/PLAITO. It is part of a collaborative research program between the University of Ottawa’s Nanda laboratory and the SnT centre at the University of Luxembourg.
Conflict of Interest The authors declared that they have no conflict of interest.
Ethical Approval Not applicable.
Informed Consent Not applicable.
Clinical Trial Number Not applicable.

### v6

@@ 6 Related Work
Requirements traceability (RT) has been extensively studied in RE mucha2024systematic ; li2023applications ; tufail2017systematic ; wang2018requirements ; ramesh1998factors . Existing work employs a range of technologies, from traditional methods such as Information Retrieval (IR) and statistical models to more advanced approaches such as Machine Learning (ML) and Deep Learning (DL). Early works borrowed IR techniques such as Vector Space Models (VSM), Latent Dirichlet Allocation (LDA), to find trace links between software artifacts via text relevancy gao2022using ; guo2017tackling ; kuang2017analyzing ; nishikawa2015recovering ; panichella2013and ; Mahmoud2013 ; Wang20193 ; sun2017frlink ; capobianco2013improving ; chhabra2017filtering ; bavota2014enhancing ; shao2013improved ; Wang20191 ; Li2020 ; moran2020improving . More advanced techniques have been introduced using ML Rasiman2022 ; Bella2019 ; mills2019tracing ; Bella2018 ; Mills2018 ; Rath2018 ; Zhao2018 ; Mills2017 ; Mills20172 ; Falessi2017 ; Hayes2015 ; le2015rclinker ; li2015recovering and DL kenton2019bert ; Zhang2021 ; Wang20192 ; Guo:17 ; Chen2019 ; Alhoshan2019 ; Alhoshan2018 ; Li2018 ; zhao2020extended ; Sultanov2013 , employing various algorithms — from classifiers like SVM, random forest, and decision trees to more sophisticated language models like BERT devlin-etal-2019-bert to find trace links. In recent years, with the emergence of LLMs, researchers have leveraged pre-trained knowledge via prompt engineering to identify trace links among software artifacts hassine2024llm ; rodriguez2023prompts ; vogelsang2025impact . Hassine hassine2024llm proposed an LLM-based technique that uses zero-shot learning on GPT3.5 to find trace links between requirements and goals in Goal-oriented Language (GRL) models. Moreover, Rodriguez et al. rodriguez2023prompts proposed an approach that integrates zero-shot prompting with reasoning to improve performance on the Traceability Link Recovery (TLR) problem across diverse software artifacts. They have shown that a prompt that performs well with one model or dataset may not yield optimal results with another, underscoring the need to tailor prompts to the specific context. Recent studies have explored the use of prompting techniques for traceability within requirements fuchss2025lissa ; ronanki2024requirements ; fuchss2025beyond ; hey2025requirements . Hey et al. hey2025requirements used Retrieval-Augmented Generation (RAG) with LLMs for inter-requirements traceability, where the model first retrieves relevant candidate requirements. Then, it identifies trace links among the retrieved candidates using the KISS prompt, which asks the model whether a trace link exists between a given pair of artifacts. Fuchß fuchss2025beyond et al. proposed an ensemble of LLM-based prompts for candidate filtering, where multiple LLM models are iteratively used to re-rank or filter the initial set of retrieved candidates. The prompt template used in this work follows the same design as that proposed by Hey et al. hey2025requirements . Additionally, Ronanki et al. ronanki2024requirements developed and evaluated five distinct interactive prompting patterns, each tuned to a specific dataset.
@@
Legal requirements traceability has only been investigated to a limited extent in the literature. Cleland-Huang et al. cleland:2010 propose a probabilistic approach that identifies trace links between requirements and the HIPAA regulation by computing probabilities based on the detection of indicator terms for regulations that correspond to requirements. The authors further propose extending the indicator terms with more domain-specific terms retrieved from the web. In follow-up work, Gibiec et al. Gibiec:2010 further investigate web mining. Guo et al. Guo:17 extend the previous two papers to improve the terminology gap problem, i.e., the mismatch between terms in requirements and regulations. The authors investigate methods based on classification, ontologies, and web mining and evaluate them on HIPAA.
While previous research has made significant strides in requirements traceability using traditional IR methods and ML/DL techniques, these approaches exhibit notable limitations in addressing the complexities of the LRT task. Most notably, existing methods struggle with the terminology gap between regulations and technical requirements, do not generalize well across regulations, and lack adaptability to multi-domain applications. Additionally, studies proposing prompt-based approaches have not fully leveraged the capabilities of prompt engineering, providing only detailed instructions and limited context. Their approaches primarily rely on pre- and post-processing strategies hey2025requirements ; fuchss2025beyond , which, in some cases, query multiple large language models. In comparison to the above works, we empirically evaluate two automated LRT approaches: (1) a classifier-based solution leveraging sentence transformers and (2) a generative LLM-based solution guided by structured prompt engineering. By exploring these methods across two distinct regulations, HIPAA and GDPR, we advance the understanding of how modern NLP techniques can be adapted to meet the challenges of LRT. We also shed light on the possibilities, or lack thereof, of transfer learning across regulations. To the best of our knowledge, we are also among the first to identify the strengths and limitations of LLMs in this context. Further, larger studies involving human experts are required to establish the benefits of LLMs for LRT.
@@ 7 Conclusion
This study presents a comparative evaluation of two approaches to Legal Requirements Traceability (LRT): a classifier-based method, $\mathcal{K}$ashif, leveraging Sentence Transformers, and a generative LLM-based method, Rice_LRT, designed using a structured prompt engineering framework. Moreover, we compare $\mathcal{K}$ashif against seven distinct baselines—ranging from traditional IR techniques to recent deep learning and transformer-based models (LSI, LDA, GloVe, TraceBERT, RoBERTa, and LLaMa2)—to comprehensively demonstrate its capabilities. Our results demonstrate that $\mathcal{K}$ashif provides significant improvements over the baselines in terms of F2 score, achieving a F2 score of $\approx$63% on HIPAA data (23 pp more than the best baseline). On the other hand, $\mathcal{K}$ashif’s performance deteriorates on more complex datasets such as GDPR, yielding only 15% recall. However, this also shows that $\mathcal{K}$ashif, despite being tuned on a similar but less complex dataset, achieves performance comparable to a simple prompt (answering yes/no) on GPT-4o, which does not fully exploit its capabilities.
Conversely, the Rice_LRT approach, built on generative LLMs, outperformed $\mathcal{K}$ashif on GDPR data, achieving an F2 score of 61% and reducing manual effort for traceability by enabling analysts to vet only a fraction of trace links. Moreover, Rice_LRT outperformed a simpler prompt, increasing TPs while reducing FPs. These findings suggest that generative LLMs and carefully designed prompts offer a promising approach to automating LRT tasks in complex legal domains. However, the approach has its challenges, including the occurrence of false positives that require further investigation. In addition to evaluating current state-of-the-art methods, this work highlights critical challenges, including terminology gaps between requirements and provisions By addressing these challenges, our study underscores the importance of tailoring solutions to the nuances of legal and regulatory contexts.
@@
Acknowledgement. This work is supported by the H2020 COSMOS European project, grant agreement No. 957254, NSERC of Canada under the Discovery and CRC programs, the Research Ireland grant 13/RC/2094-2, the Luxembourg National Research Fund under grant numbers C23/IS/17958091/PLAITO and NCER22/IS/16570468/NCER-FT. It is part of a collaborative research program between the University of Ottawa’s Nanda laboratory and the SnT centre at the University of Luxembourg.
Declaration
Conflict of Interest. The authors declared that they have no conflict of interest.